In ten years Mark will be 8 times as old as he was four years ago. Find his present age
Answer by macston(5194)

In ten years Mark will be 8 times as old as he was four years ago. Find his present age
M=Mark's age
M+10=8(M-4)
M+10=8M-32 Add -M+32 to each side.
42=7M Divide each side by 7.
6=M ANSWER: Mark is 6 years old.